# Knowledge Transfer from High-Resource to Low-Resource Programming Languages for Code LLMs

### v5

@@ 6. Discussion
We have shown that MultiPL-T is an effective and efficient method for generating semi-synthetic training data for low-resource programming languages. In this section we discuss the implications of extending MultiPL-T in various ways.
@@
We hope it is clear to the reader that the MultiPL-T approach is straightforward to generalize to more programming languages. The language-specific work involves 1) translating comments and function signatures into an appropriate prompt, and 2) writing a compiler that can translate simple assertions from the source to target. MultiPL-E already supports both these steps for 20+ programming languages, several of which are low-resource, including D, Bash, MATLAB, Haskell, and Perl. So, generating fine-tuning sets for these languages may just be a matter of running the MultiPL-T pipeline for a few days on GPUs.
@@
Although this paper fine-tunes the Code Llama and StarCoder family of Code LLMs, our datasets could also be used to fine-tune other LLMs. We use StarCoderBase-15B to generate data, but it would be straightforward to swap in some other model as well.
@@
Although we have argued that self-instruct is unlikely to succeed on a low-resource language, self-instruct and MultiPL-T could be composed together in a natural way: one could generate a high-quality dataset of instructions in a high-resource language, and then use MultiPL-T to translate them to a low-resource language. Given the effectiveness of WizardCoder (Luo et al., 2023) at Python, this composition seems likely to succeed.
@@ 7. Related Work
A number of projects use language models to translate code between programming languages and test that the generated translations are correct by compiling working unit tests from one language to another. TransCoder-ST (Roziere et al., 2021) and CMTrans (Xie et al., 2024) uses these techniques to generate training data for a code translation model between Java, Python, and C++, whereas MultiPL-E (Cassano et al., 2023), MBXP (Athiwaratkun et al., 2022), and BabelCode (Orlanski et al., 2023) translate Code LLM benchmarks from Python to 10+ programming languages. A distinguishing feature of MultiPL-T is that it employs an off-the-shelf pretrained Code LLM (StarCoder) to both generate test cases and translate code to low-resource languages. When the aforementioned papers were written, the best open Code LLMs were far less capable than StarCoder: they were trained on fewer programming languages using far less training data, and they were an order of magnitude smaller. Thus we believe the MultiPL-T approach would have likely failed. Although capable closed models were available, they were either rate limited or prohibitively expensive for the scale of data generation that MultiPL-T needs. For example, Cassano et al. (2023) report that they used a commercial model during a free beta period, but it would have cost \$37,000 with equivalent released models. MultiPL-T’s data generation requires an order of magnitude more queries, making it prohibitively expensive to use with commercial models at 2023 prices.
@@
To get an LLM to perform a desired task, the user must prompt it in the right way. There are several techniques for instruction tuning LLMs to better follow natural, human-written instructions. One approach uses human annotators to write sample instructions and give feedback on a large number of model generations (Ouyang et al., 2022), but this is expensive and requires significant resources. A cheaper approach is to have a capable LLM self-instruct to generate instructions from a relatively small set of human-written seed instructions (Wang et al., 2023). Evol-Instruct uses an LLM to create variations of instructions (Luo et al., 2023). These techniques have been used to create datasets for instruction-tuning Code LLMs (Chaudhary, 2023; Muennighoff et al., 2023; Luo et al., 2023). These datasets focus on high-resource languages, and, as we show in § 3.2, they are unlikely to succeed for low-resource languages.
Training on high quality data
Training on high-quality data is an effective way to reduce both the size of an LLM and the volume of training data needed, while maintaining performance. Gunasekar et al. (2023) achieve high HumanEval scores on a small model with a modest amount of “textbook quality” training data. This includes both natural and synthetic data generated by a more capable model. Their work targets Python, and we argue in § 3 that the approach is less likely to succeed on low-resource languages.
@@
Models fine-tuned on the output of proprietary Code LLMs
Despite the fact that many proprietary models forbid using their generated data to train other models, there are models that do so. The Phind and WizardCoder (Luo et al., 2023) projects have fine-tuned both StarCoder and Code Llama models on data generated by GPT-3.5 and GPT-4. Phind also uses additional proprietary data that is orders of magnitude larger than the MultiPL-T training data. These models outperform our fine-tuned Code Llama on R, and are close on the other programming languages.
However, we do not believe these comparisons are very meaningful. One could fine-tune these models further on the MultiPL-T data we already have, or even use them (or GPT-4) to translate Python to low-resource languages. But, we would not be able to compare to the natural baseline of training longer on existing data, which for these models is the proprietary OpenAI training set. The strength of the MultiPL-T approach is not that it does better in an absolute sense, but that fine-tuning with MultiPL-T data is better than training longer on existing data. We show that this holds for StarCoderBase and its training set, and can only do so because both are open.
@@ 8. Conclusion
In the last few years, Code LLMs have rapidly made their way into more and more programming tools. However, the quality and reliability of Code LLMs is highly language-dependent: they can be remarkable on high-resource programming languages, but are far less impressive at working with low-resource languages. It is possible that in the near future, a large number of programmers will expect LLM-based technology to work, just as many programmers today expect syntax highlighting, continuous analysis, or type-based completion in their programming environments. We hope that MultiPL-T—a methodology for generating large-scale, high-quality fine-tuning datasets in low-resource languages—will help low-resource languages compete in a world where many developer tools rely on Code LLMs.
@@
The MultiPL-T fine-tuning data (and code) are also open: they are constructed from the StarCoder training data (The Stack) and augmented by StarCoder itself. We deliberately do not use a more capable proprietary model to fine-tune an open model. This allows us to demonstrate that fine-tuning on MultiPL-T data is more effective and efficient than training longer on existing data. We evaluate MultiPL-T in several other ways, including a new adversarial benchmark and a qualitative evaluation of coding style. When compared to open models that are not fine-tuned on proprietary data, MultiPL-T achieves state-of-the-art results for Julia, Lua, OCaml, R, and Racket.
Appendix A Human-Written Prompt for Self Instruct
We use the following five functions to start a round of self-instruct for Racket. We generate results at temperature 0.8.
;; watching-you?: Number -> Boolean
;; Sees if the number has a "00" in it
(define (watching-you? num)
    (string-contains? (number->string num) "00"))
;; add-odds: [List-of Numbers] -> Number
;; Adds all the odd numbers in a list
(define (add-odds lon)
    (foldr + 0 (filter odd? lon)))
;; repeat: String Number -> [List-of Strings]
;; Repeats a string num amount of times
(define (repeat str num)
    (if (= num 0) ""
    (string-append str (repeat str (- num 1)))))
;; sum-remainders: [List-of Numbers] Number -> Number
;; Sums the remainder of the numbers in the list when divided by num
(define (sum-remainders lon num)
    (foldr + 0 (map (lambda (n) (remainder n num)) lon)))
;; is-palindrome?: Number -> Boolean
;; Checks if a number is a palindrome
(define (is-palindrome? num)
    (equal? (number->string num)
            (list->string
                (reverse (string->list (number->string num))))))
Appendix B More Details on Evaluation
Language StarCoderBase-1B StarCoderBase-15B CodeLlama-34B
Epoch
OCaml 3 4 2
Racket 6 8 4
Lua 6 4 2
R 6 2 4
Julia 4 6 4
Table 7. Epochs chosen for each language and model fine-tuned as described in § 5.
The results in table 2 and table 4 report the best performance of the fine-tuned models on MultiPL-E, which we find at the epochs highlighted in Table 7.
B.1. Evaluation On MultiPL-MBPP
Language StarCoderBase-1B StarCoderBase-15B CodeLlama-34B
Base Fine-tuned Base Fine-tuned Base Fine-tuned
OCaml 3.7 20.2 17.2 28.9 28.9 35.8
Racket 4.7 19.3 21.1 30.5 29.8 39.7
Lua 18.7 26.6 35.2 40.3 42.8 47.4
Table 8. MultiPL-MBPP pass@1 scores for 1B, 15B, and 34B parameter models fine-tuned on MultiPL-T data.
The main body of this article focuses on the MultiPL-E translation of HumanEval, but MultiPL-E also translates MBPP (Austin et al., 2021), which is a program synthesis benchmark comprised of 427 crowdsourced Python problems. This benchmark is similar to MultiPL-E, but more than twice the number of problems and is typically considered to be less challenging. We further evaluate our fine-tuned models for OCaml, Racket, and Lua MultiPL-MBPP.
Appendix C Pass Rate with Adversarial Benchmark
Comparison of StarCoderBase-15B models’ performance when they were just trained on the Stack versus after being fine-tuned on MultiPl-T dataset.

### v6

@@ 6. Discussion
We have shown that MultiPL-T is an effective and efficient method for generating semi-synthetic training data for low-resource programming languages. In this section, we discuss the implications of extending MultiPL-T in various ways.
@@
We hope it is clear to the reader that the MultiPL-T approach is straightforward to generalize to more programming languages. The language-specific work involves 1) translating comments and function signatures into an appropriate prompt, and 2) writing a compiler that can translate simple assertions from the source to the target. MultiPL-E already supports both these steps for 20+ programming languages, several of which are low-resource, including D, Bash, MATLAB, Haskell, and Perl. So, generating fine-tuning sets for these languages may just be a matter of running the MultiPL-T pipeline for a few days on GPUs.
Some combinations of source and target languages may be more effective than others. For example, consider building semi-synthetic training data for Rust. We could directly leverage the datasets in this paper and utilize a Code LLM to translate Python to Rust. However, we speculate that it would be more effective to modify MultiPL-T to support a source language closer to Rust, such as C++.
@@
Although this paper focuses on fine-tunes of the Code Llama and StarCoder family of Code LLMs, our datasets could also be used to fine-tune other LLMs. We fine-tune DeepSeek Coder and StarCoder2 on Racket, and we expect our results will generalize to the other languages.
Regenerating MultiPL-T data
The MultiPL-T pipeline in this paper uses StarCoderBase-15B. It should be clear that using a more capable model would improve conversion rates, and thus produce better results. A proprietary model such as GPT-4 would likely produce the best results, but doing so would violate its terms of use (OpenAI, 2023b). An unusual result in this paper is that MultiPL-T can use a weaker model (StarCoderBase-15B) to improve much better models (CodeLlama-70B, DeepSeekCoder-34B, and StarCoder2-15B) on low-resource languages.
@@
Although we have argued that self-instruct is unlikely to succeed on a low-resource language, self-instruct and MultiPL-T could be composed together in a natural way: one could generate a high-quality dataset of instructions in a high-resource language, and then use MultiPL-T to translate them to a low-resource language. Given the effectiveness of WizardCoder (Luo et al., 2024) and Magicoder (Wei et al., 2024) at Python, this composition seems likely to succeed.
Other kinds of benchmarks
A lot of effort has been put into evaluating the Python programming abilities of LLMs, but there are far fewer benchmarks for low-resource languages. For example, there are Python benchmarks for specialized tasks, such as data science (Lai et al., 2023), and with prompts authored by specific populations, such as beginning programmers (Babe et al., 2024). We need to develop these kinds of benchmarks for low-resource languages to truly understand the capabilities and limitations of Code LLMs.
@@ 7. Related Work
A number of projects use language models to translate code between programming languages and test that the generated translations are correct by compiling working unit tests from one language to another. TransCoder-ST (Roziere et al., 2021) and CMTrans (Xie et al., 2024) use these techniques to generate training data for a code translation model between Java, Python, and C++, whereas MultiPL-E (Cassano et al., 2023), MBXP (Athiwaratkun et al., 2022), and BabelCode (Orlanski et al., 2023) translate Code LLM benchmarks from Python to 10+ programming languages. A distinguishing feature of MultiPL-T is that it employs an off-the-shelf pretrained Code LLM (StarCoder) to both generate test cases and translate code to low-resource languages. When the aforementioned papers were written, the best open Code LLMs were far less capable than StarCoder: they were trained on fewer programming languages using far less training data, and they were an order of magnitude smaller. Thus we believe the MultiPL-T approach would have likely failed. Although capable closed models were available, they were either rate-limited or prohibitively expensive for the scale of data generation that MultiPL-T needs. For example, Cassano et al. (2023) report that they used a commercial model during a free beta period, but it would have cost \$37,000 with equivalent released models. MultiPL-T’s data generation requires an order of magnitude more queries, making it prohibitively expensive to use with commercial models at 2023 prices.
Multi-lingual language models of code
Training models on a dataset of code written in programming languages similar to the target language, known as transfer learning, is a commonly studied method for improving model performance on a specific programming language (Ahmed and Devanbu, 2022; Baltaji et al., 2024; Katzy et al., 2023). Chen et al. (2022) explore this method in the context of low-resource languages by pretraining and fine-tuning small encoder-only Transformers with different programming languages. They then evaluate their performance on code search and summarization tasks for Ruby, a language with a dataset 10x smaller than that of Python. Their findings suggest that multilingual pretraining can improve performance at generating code in low-resource languages. However, this strategy may falter with languages that significantly differ in syntax or semantics, such as Racket. For example, Baltaji et al. (2024) employ a similar methodology to train models on 41 different programming language pairs, noting that Scheme, the predecessor of Racket, benefits less from transfer learning compared to other languages. In such scenarios, data generation strategies like MultiPL-T are deemed more likely to achieve success. Moreover, these studies predominantly rely on the BLEU score and other syntax-based metrics for evaluating correctness, which have been criticized for their inadequacy in code generation tasks (Chen et al., 2021). In our work, we utilize test-based validation of both our training data and evaluation metrics to ensure the correctness of generated code, which is a more reliable measure of model performance.
Our work begins with the observation that Code LLM performance can vary significantly by language, and this has been observed repeated in prior work. Most closely related to MultiPL-T are benchmarks for the natural language to code task (Cassano et al., 2023; Orlanski et al., 2023; Athiwaratkun et al., 2022). However, similar trends occur when Code LLMs are used for other tasks, such as fuzzing and translation (Xia et al., 2024; Pan et al., 2024).
@@
To get an LLM to perform a desired task, the user must prompt it in the right way. There are several techniques for instruction tuning LLMs to better follow natural, human-written instructions. One approach uses human annotators to write sample instructions and give feedback on a large number of model generations (Ouyang et al., 2022), but this is expensive and requires significant resources. A cheaper approach is to have a capable LLM self-instruct to generate instructions from a relatively small set of human-written seed instructions (Wang et al., 2023). Evol-Instruct uses an LLM to create variations of instructions (Luo et al., 2024). These techniques have been used to create datasets for instruction-tuning Code LLMs (Chaudhary, 2023; Muennighoff et al., 2024; Luo et al., 2024; Wei et al., 2024). These datasets focus on high-resource languages, and, as we show in § 3.2, they are unlikely to succeed for low-resource languages.
Training on high-quality data
Training on high-quality data is an effective way to reduce both the size of an LLM and the volume of training data needed, while maintaining performance. Gunasekar et al. (2023) achieves high HumanEval scores on a small model with a modest amount of “textbook quality” training data. This includes both natural and synthetic data generated by a more capable model. Their work targets Python, and we argue in § 3 that the approach is less likely to succeed with low-resource languages.
@@
Even though, many proprietary models forbid using their generated data to train other models, there are models that do so. These models are distillations of much larger proprietary models, and it is not possible to compare them to the natural baseline, which is training longer on existing data, which for these models is the proprietary OpenAI training set. The strength of the MultiPL-T approach is not that it does better in an absolute sense, but that fine-tuning with MultiPL-T data is better than training longer on existing data. We show that this holds for StarCoderBase and its training set, and can only do so because both are open.
@@ 8. Conclusion
In the last few years, Code LLMs have rapidly made their way into more and more programming tools. However, the quality and reliability of Code LLMs is highly language-dependent: they can be remarkable on high-resource programming languages, but are far less impressive at working with low-resource languages (Cassano et al., 2023). It is possible that in the near future, a large number of programmers will expect LLM-based technology to just work, just as many programmers today expect syntax highlighting, continuous analysis, or type-based completion in their programming environments. We hope that MultiPL-T—a methodology for generating large-scale, high-quality fine-tuning datasets in low-resource languages—will help low-resource languages compete in a world where many developer tools rely on Code LLMs.
@@
The MultiPL-T fine-tuning data (and code) are also open: they are constructed from the StarCoder training data (The Stack) and augmented by StarCoder itself. We deliberately do not use a more capable proprietary model to fine-tune an open model. This allows us to demonstrate that fine-tuning on MultiPL-T data is more effective and efficient than training longer on existing data. We evaluate MultiPL-T in several other ways, including a new benchmark that exercises in-context learning and a qualitative evaluation of coding style. When compared to other fine-tunes of the same base model, MultiPL-T achieves state-of-the-art results for Julia, Lua, OCaml, R, and Racket.
Data-Availability Statement
All code, data, and models from this paper are available with open licenses. Code: available on GitHub and archived on Zenodo; Datasets: available on Hugging Face; and Models: available on Hugging Face. In particular:
  1. (1)
    A guide to reproduce the results in this article is available at doi.org/10.5281/zenodo.12453932.
  2. (2)
    The MultiPL-T datasets are available at doi.org/10.57967/hf/2941 with links to several models fine-tuned on these datasets.
Acknowledgements
We thank Loubna Ben Allal, Harm de Vries, Joydeep Biswas, Matthias Felleisen, Shriram Krishnamurthi, and Leandro von Werra for helpful conversations. Thanks to Leandro von Werra for including the MultiPL-T datasets in the StarCoder2 training corpus. We also thank the OOPSLA reviewers for their feedback. We especially thank the artifact evaluation committee for their patience reviewing our complex artifact, as well as the AEC chairs (Guillaume Baudart and Sankha Narayan Guria) for facilitating review. This work used computing resources provided by Northeastern Research Computing, New England Research Cloud, and Joydeep Biswas (UT Austin). Federico Cassano was affiliated with Roblox for most of his work on this paper. This work is partially supported by the National Science Foundation (SES-2326173, SES-2326174, and SES-2326175).